Jatropha seeds have good potential for biodiesel production. Oil extraction and oil refining are essential steps before biodiesel production. During the oil extraction of 200 g seeds, it provides 110 g of pressed meat with 3% losses. If the oil refining showed the output below. How many milliliters (mL) of gums, wax, soap and oxidation products produced. The density of the Jatropha seeds oil is 880 kg/m3. [1 L = 1000 mL] Oxidation products %, 9 Wax %, 13 Soap %, 6 Gums%, 4 Refined Oil %68
Answer #1

mass of seeds = 200 g

mass of pressed meat = 110 g

total mass of oil = mass of seeds - mass of pressed meat

total mass of oil = (200 g) - (110 g)

total mass of oil = 90 g

mass of oil recovered = 97% * total mass of oil

mass of oil recovered = (97/100) * (90 g)

mass of oil recovered = 87.3 g

volume of oil recovered = (mass of oil recovered) / (density of oil)

volume of oil recovered = (87.3 g) / (0.880 g/mL)

volume of oil recovered = 99.2 mL

Volume of oxidation products = (% oxidation products) * (volume of oil recovered)

Volume of oxidation products = (9 %) * (99.2 mL)

Volume of oxidation products = (9/100) * (99.2 mL)

Volume of oxidation products = 8.928 mL

Volume of wax = (% wax) * (volume of oil recovered)

Volume of wax = (13 %) * (99.2 mL)

Volume of wax = (13/100) * (99.2 mL)

Volume of wax = 12.896 mL

Volume of soap = (% soap) * (volume of oil recovered)

Volume of soap = (6 %) * (99.2 mL)

Volume of soap = (6/100) * (99.2 mL)

Volume of soap = 5.952 mL

Volume of gums = (% gums) * (volume of oil recovered)

Volume of gums = (4 %) * (99.2 mL)

Volume of gums = (4/100) * (99.2 mL)

Volume of gums = 3.968 mL

Volume of refined oil = (% refined oil) * (volume of oil recovered)

Volume of refined oil = (68 %) * (99.2 mL)

Volume of refined oil = (68/100) * (99.2 mL)

Volume of refined oil = 67.456 mL